Orwell building supplies' last dividend was $1.75. Its dividend growth rate is expected to be constant at 37.00% for 2 years, after which dividends are expected to grow at a rate of 6% forever. Its required return (rs) is 12%. What is the best estimate of the current stock price?

Select the correct answer.

a. $49.84
b. $51.02
c. $53.38
d. $54.56
e. $52.20
Answer #1

Please find below the solution ... Please do provide rating...

Correct answer is option b. $51.02

Current price is present value of future cash flow/Dividend
Year - 1 dividend = 1.75*137%          2.40
Year -2 dividend = 2.4*137%          3.28
At the end of year-2 terminal value.
Terminal cash flow = 3rd year dividend /(required rate - Growth rate)
=3.28*1.06/(12%-6%)
      58.03
Year Dividend Horizon value Total cash flow PVIF @ 12% present value
1          2.40          2.40     0.8929          2.14
2          3.28       58.03       61.31     0.7972       48.88
      51.02